A 5L gas is under a pressure 2 kPa at a temperature of 15C. What will the temperature be if the volume changes to 15L


1
Expert's answer
2021-05-27T07:34:44-0400

V=vRT/p

R(constant gas temperature)=8.31J

T=V*p/vR

v=15l/22.4l/mol=0.67mol

T=15l*2kPa/0.67mol*8.31J=5.4C